OpenStudy (igster1112):

One cantaloupe costs $2.75 How many cantaloupes can you buy for $30.25?

OpenStudy (mathstudent55):

This is a division problem.

OpenStudy (mathstudent55):

You want to know how many times 2.75 goes into 30.25

OpenStudy (studygurl14):

\(\Large \frac{$30.25}{$2.75}=\tt the~answer\) Round down to a whole number, if applicable.

OpenStudy (igster1112):

I have to solve it as a proportion.

OpenStudy (studygurl14):

Proportion: \(\Large\frac{2.75}{1}=\frac{30.25}{c}\) Solve for c

OpenStudy (mathstudent55):



OpenStudy (igster1112):

Oh ok

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

When you cross multiply, you do two products that cross each other (as the arrows in the drawing above show), and you set the products equal. Then you solve for the variable.

OpenStudy (igster1112):

Oh ok thank you

OpenStudy (mathstudent55):

You're welcome.